I have sent this before and I've checked the solutions on the website on this topic but I still can't find the answers. Perhaps, some of tutors can give me a hand.
Question
Two towns P and Q are 35 km apart. Peter starts cycling from P towards Q at 12 p.m. at 20 km/h until he is 16 km from P, when he changes speed so that he arrives at Q at 2 p.m. John leaves Q at 12:30 p.m. and cycles to P at a constant speed 26 km/h, find
a)the time when Peter and John meet.
b)the distance that they will meet
Thank you.

Answer by venugopalramana(3286)

THE NUMBERS GIVEN IN THIS PROBLEM ALL LEAD TO FRACTIONS INVOLVING HRS,MTS AND SECS.ARE YOU SURE YOU GOT THE NUMBERS RIGHT?IS DISTANCE BETWEN P AND Q =34 KM OR 35 KM?
Two towns P and Q are 35 km apart. Peter starts cycling from P towards Q at 12 p.m. at 20 km/h until he is 16 km from P, when he changes speed so that he arrives at Q at 2 p.m. John leaves Q at 12:30 p.m. and cycles to P at a constant speed 26 km/h, find
a)the time when Peter and John meet.
b)the distance that they will meet
PETER TRAVELS STARTING FROM P AT 12=00 PM GOING AT 20 KMPH FOR 16 KM
TIME TAKEN =16/20=0.8 HRS=0.8*60=48 MTS….THAT IS AT 12 =48
BALANCE DISTANCE TO Q =35-16=19 KM.
TIME TAKEN TO REACH =2=00 PM THAT IS 14=00 HRS
TRAVEL TIME IS 14=00 - 12 =48 =1 HR12 MTS. = 1 +12/60 HRS =6/5 HRS.
PETERS SPEED DURING THIS TRAVEL =19/(6/5)=5*19/6=95/6 KMPH
JOHN LEAVES Q AT 12=30 PM.
TILL 12=48 HE TRAVELS FOR 12=48 -12=00=18 MTS=18/60 HRS =0.3 HRS.
DISTANCE TRAVELLED BY JOHN DURING THIS PERIOD AT 26 KMPH=26*0.3=7.8 KM
HENCE DISTANCE OF SEPERATION BETWEEN PETER AND JOHN AT 12 = 48 IS (35-16)-7.8=11.2 KM
THEIR RELATIVE SPEED WHILE TRAVELLING IN OPPOSITE DIRECTIONS =95/6 +26=(95+156)/6=251/6 KMPH
TIME TAKEN TO MEET =11.2/(251/6)=67.2/251 HRS = 16.06 MTS
HENCE PETER AND JOHN MEET AT 12=48 + 16.06 MTS..THAT IS AT 13 HRS.04 MTS.36 SEC
THE DISTANCE AT WHICH THEY MEET =16.06*95/6=4.24 KM THAT IS AT 16+4.24=18.24 KM FROM P.